How long does it take light to travel from the Sun to the Earth? The Sun’s nearest neighbor among the stars, a dim little star called Proxima Centauri, is at a distance of 1.295 parsecs from us. How long does it take light to travel from Proxima Centauri to the Earth?
Answer #1

We know that

time = distance/Speed

Speed of light = 3*10^8 m/sec

distance between Sun and earth = 1.496*10^11 m

time = 1.496*10^11 m/(3*10^8 m/sec) = 498.67 sec

time = (498.67 sec)*(1 min/60 sec) = 8.31 min

Part B.

We know that

time = distance/Velocity

Speed of light = 3*10^8 m/sec

distance between proxima centauri and earth = 1.295 pc

Since 1 pc = 1 parsec = 3.086*10^16 m

So,

time = 1.295*3.086*10^16 m/(3*10^8 m/sec)

time = 133212333 sec

Now we know that 1 yr = 3.154*10^7 sec

So,

time = (133212333 sec)*(1 yr/(3.154*10^7 sec)) = 4.22 yr

Please Upvote.